# Visual-Inertial SLAM as Simple as A, B, VINS

### v2

@@ III-C Local Mapping
Every so often the current frame is selected to be a keyframe. A new keyframe is made when the percent of features visible in the current frame compared to the most recent keyframe falls below a threshold. Thus if the camera is always looking at the same thing, no new keyframe will be made. Keyframes are pushed onto a local mapping window, but first a monocular depth network (MiDaS [23, 36]) is run on the raw image  [^2] All networks are run on the raw instead of rectified image in order to benefit from the full available FoV. to obtain an affine-invariant inverse depth map which is up to a scale and bias parameter $a$ and $b$, respectively. The scale and bias, $a$ and $b$, as well as some other terms to correct the depth are estimated in the place of each feature position separately, which we call the AB feature representation.
@@ III-C1 AB Features
The watershed algorithm is then used on the edge image to create what we call a code mask, which partitions the image into $C$ different regions. Using the code mask, for each keyframe a $C$-dimensional code vector $\mathbf{c}$ is estimated to push and pull different regions of the depth map in order to correct it using multi-view information. Given a code mask element $m$ for $d_{inv}$ we can write
@@
After $\mathbf{c}$ of course comes $\mathbf{d}$. We estimate a 4-dimensional vector $\mathbf{d}$ in order to make up for the fact that we use normalized bearings in AB-VINS, and run the MiDaS network on raw distorted images, while it was trained on a simple pinhole model (because the network predicts inverse depths that are too large near the edge of fisheye images). For $\mathbf{d}$ we write
$r=||\mathbf{u}_{n}||~{}~{}~{}~{}$ $\theta=\mathrm{atan2}(r,1)$
$\theta_{d}=\theta(1+d[0]\theta^{2}+d[1]\theta^{4}+$ $d[2]\theta^{6}+d[3]\theta^{8})~{}~{}~{}~{}d_{c}=\theta_{d}/r$
@@
where $\mathbf{u}_{n}$ is the ideal image coordinate, which is of course a function $\mathbf{u}_{n}=\Pi^{-1}(\bm{\zeta},\mathbf{u})$ of the camera intrinsics $\bm{\zeta}$ and raw image coordinate $\mathbf{u}$. Thus $\mathbf{d}$ makes the inverse depth smaller closer to the edges of the image. Since Eq. 8 is nearly identical to the Kannala-Brandt [38] fisheye lens distorting equations, we call the process of $\mathbf{d}$ fisheye monocular depth. The estimation of $a$ and $b$, with the optional $\mathbf{c}$ and $\mathbf{d}$ terms, to represent feature positions is called the AB feature representation.
@@
${}^{C}\mathbf{p}_{f}$ $=\frac{\begin{bmatrix}\mathbf{u}_{n}^{\top}&1\end{bmatrix}^{\top}}{z_{inv}\lVert\begin{bmatrix}\mathbf{u}_{n}^{\top}&1\end{bmatrix}^{\top}\rVert}$ (9)
@@ III-C2 Depth Map Registration
Before being pushed onto the local mapping window, the depth map is registered to the current local map – linearly estimating $a$ and $b$ using the linear system presented in [23] with the sparse points projected from the local map (the same points used in tracking to static map) as the reference metric inverse depths. Note that we wrap the linear registration in a RANSAC loop to improve the robustness. Linear registration is followed by a nonlinear optimization to jointly estimate $a$, $b$, $\mathbf{c}$, and $\mathbf{d}$ while holding the local map fixed – using the same 3D points as in tracking to static map for both linear and nonlinear registration. The main cost in this optimization is the depth consistency cost, which is similar to the depth consistency costs in [21, 14] but defined in the inverse depth space in order for the cost to focus on closer objects rather than farther ones. The $\mathbf{c}$ and $\mathbf{d}$ terms are initialized to zero, and a zero prior is placed on $\mathbf{c}$ and $\mathbf{d}$ in order to prevent the terms from blowing up, which is similar to the zero-code prior in [21, 14]. In particular, the nonlinear depth map registration attempts to minimize the cost function $\mathcal{C}_{dr}$ defined as
@@
$\hat{z}_{inv_{\ell k}}$ is the estimated metric inverse depth of the $\ell$th feature at time $t_{k}$, and $h_{ds}$ projects the 3D point ${}^{A}\mathbf{p}_{\ell}$ into frame $k$ and extracts the inverse depth. $\sigma_{d}$ is the manually-tuned noise value for the depth consistency residual. We also have the prior residual
@@ III-C3 Local Mapping Optimization
After registration, as shown in Fig. 5 we perform a sliding-window optimization in the local mapper, with the oldest pose, velocity, and biases fixed. Online calibration of the camera intrinsics and extrinsics is also performed in this optimization. The default sliding window size of our system is 9.
All of the poses are defined in the anchored $\{A\}$ frame, which is the gravity-aligned frame centered around the oldest keyframe’s frame of reference, and is changed after the oldest keyframe is popped from the window. This is similar to the robocentric VIO formulation in [11], but [11] defines the robocentric frame to be the oldest frame of reference in the sliding window, which requires actually estimating the gravity direction. On the other hand, our anchored frame is gravity-aligned – only yaw-rotated when the oldest keyframe is popped off the window – and does not required estimating gravity. While the global pose is not tracked in the local mapper, AB-VINS still calculates it for system output by using the most recent global pose in the memory tree to transform poses in the $\{A\}$ frame into the global  [^3] Note that traversal of the memory tree to calculate the global pose, which is logarithmic in the number of keyframes, is done solely in the mapping thread to keep VIO constant-time..
Besides the standard reprojection error, our local mapping optimization also includes depth consistency cost. Due to having a dense depth map for each keyframe, every bearing in the sliding window is actually its own landmark, and the landmarks simply have to work together with the oldest landmark via the depth consistency cost to keep the system stable. Once the oldest bearing for a particular feature track is older than the sliding window, it becomes a static 3D landmark ${}^{A}\mathbf{p}_{f_{\ell}}$, and replaces the AB feature for the reprojection error and depth consistency. More specifically, the local mapping optimization tries to minimize the cost function $\mathcal{C}_{lm}$, where
@@
where $\hat{z}_{inv_{\ell a}}$ is the metric inverse depth in the anchor frame for this 3D point at keyframe index $a$. Similarly, $\mathbf{r}^{r}_{\ell i}$ is defined as
@@
III-D Mapping with the Memory Tree
After the oldest keyframe is popped off of the local mapping window, it is sent to the mapper, which is responsible for finding and solving for loop closures. As shown in Fig. 2, right before entering the mapper, the keyframe receives a global descriptor from NetVLAD [39] and local descriptors from SuperPoint [40], which are typically different from the FAST corners used by VIO. At this point, in order to save RAM memory, all image-resolution dense matrices (e.g., depth map and code mask) are deleted and only sparse portions are maintained at the local descriptor locations. The global descriptor is PCA compressed to 512 dimensions, and searching for loop candidates with this descriptor is extremely efficient. The local descriptors are used to match local keypoints between loop candidates, and if enough matches are found, Horn’s method [41] wrapped in RANSAC is used to compute the relative pose for the pose graph optimization. Upon a loop closure, the loop measurements are sent to the memory tree. The memory tree is a novel data structure used to significantly speed up pose graph optimization. In the following, the details of the memory tree are presented.
III-D1 Tree Structure
As seen in Fig. 6, poses in the memory tree are defined in the parent frame within a binary search tree. Only the root’s pose is defined in the global frame. While theoretically, any pose dimension can be used in the memory tree (e.g., 6-DoF SE(3) or 3-DoF SE(2)), since VIO only drifts in 3D position and yaw we adopt the 4-DoF poses
@@ III-D1 Tree Structure
where ${}_{i}^{j}\theta_{z}$ is the yaw angle, $\mathbf{R}_{z}$ constructs a yaw-only 3D rotation matrix, and ${}^{i}\mathbf{p}_{j}$ is the gravity-aligned position. The keyframe’s timestamp is the search key for the tree. The memory tree is balanced using AVL tree rotation rules [42], which guarantee a perfectly balanced tree no matter what. Since the tree is always balanced, calculating the global pose for any keyframe can be done in $\mathcal{O}(\log(N))$ time if there are $N$ nodes in the tree by repeatedly transforming the pose into the same frame as the parent’s pose until the root node is reached. The benefit of defining the poses in this way is that optimizing a single pose also moves the entire subtree underneath the optimized pose’s node.
@@
To limit the memory usage in areas of repeated visitation, redundant nodes are pruned from the memory tree. A node is considered redundant if a certain percent of the map point visible in it’s keyframe are visible in at least three other frames, which is similar to the keyframe culling criteria of ORB-SLAM [2]. After a node is pruned, the tree is rebalanced, again using the AVL balancing rules, to ensure that the height of the tree is $\mathcal{O}(\log(N))$
III-D2 Optimizing the Memory Tree

### v3

@@ III-C Local Mapping
Every so often the current frame is selected to be a keyframe. A new keyframe is made when the percent of features visible in the current frame compared to the most recent keyframe falls below a threshold. Thus if the camera is always looking at the same thing, no new keyframe will be created. Keyframes are pushed onto a local mapping window, but first a monocular depth network (MiDaS [2, 42]) is run on the raw image to obtain an affine-invariant inverse depth map which is up to a scale and bias parameter ($a$ and $b$, respectively). All networks in AB-VINS are run on the raw distorted instead of undistorted image in order to benefit from the full available FoV. The scale and bias, $a$ and $b$, as well as some other terms to correct the depth are estimated in the place of each feature position separately, which we call the AB feature representation.
@@ III-C1 AB Features
As shown in Fig. 2, the watershed algorithm is then used on the edge image to create what we call a code mask, which partitions the image into $C$ different regions. Using the code mask, for each keyframe a $C$-dimensional code vector $\mathbf{c}$ is estimated to push and pull different regions of the depth map in order to correct it using multi-view information. Given a code mask element $m$ for $d_{inv}$ we can write
@@
After $\mathbf{c}$ of course comes $\mathbf{d}$. We estimate a 4-dimensional vector $\mathbf{d}$ in order to make up for the fact that we use normalized bearings in AB-VINS, and run the MiDaS network on raw distorted images, while it was trained on a simple pinhole model (because the network predicts inverse depths that are too large near the edge of distorted images). For $\mathbf{d}$ we write
@@
$d_{c}$ $=\theta_{d}/r$ (9)
$\theta_{d}$ $=\theta(1+d[0]\theta^{2}+d[1]\theta^{4}+d[2]\theta^{6}+d[3]\theta^{8})$ (10)
$r$ $=||\mathbf{u}_{n}||$ (11)
[…]
where $\mathbf{u}_{n}$ is the ideal image coordinate, which is a function $\mathbf{u}_{n}=\Pi^{-1}(\bm{\zeta},\mathbf{u})$ of the camera intrinsics $\bm{\zeta}$ and raw image coordinate $\mathbf{u}$. Thus $\mathbf{d}$ makes the inverse depth smaller closer to the edges of the image. Since Eq. 8 is nearly identical to the Kannala-Brandt [44] fisheye lens distorting equations, we call the process of $\mathbf{d}$ fisheye monocular depth. The estimation of $a$ and $b$, with the optional $\mathbf{c}$ and $\mathbf{d}$ terms, to represent feature positions is called the AB feature representation.
@@
${}^{C}\mathbf{p}_{f}$ $=\frac{\begin{bmatrix}\mathbf{u}_{n}^{\top}&1\end{bmatrix}^{\top}}{z_{inv}\left|\left|{\begin{bmatrix}\mathbf{u}_{n}^{\top}&1\end{bmatrix}^{\top}}\right|\right|}$ (13)
@@ III-C2 Depth Map Registration
Before being pushed onto the local mapping window, the depth map is registered to the current local map – linearly estimating $a$ and $b$ using the linear system presented in [2] with the sparse points projected from the local map as the reference metric inverse depths. Note that we wrap the linear registration in a RANSAC loop to improve the robustness. Linear registration is followed by a nonlinear optimization to jointly estimate $a$, $b$, $\mathbf{c}$, and $\mathbf{d}$ while holding the local map fixed – using the same 3D points as in tracking to static map for both linear and nonlinear registration. The main cost in this optimization is the depth consistency cost, which is similar to the depth consistency costs in [27, 20] but defined in the inverse depth space in order for the cost to focus on closer objects rather than farther ones. The $\mathbf{c}$ and $\mathbf{d}$ terms are initialized to zero, and a zero prior is placed on $\mathbf{c}$ and $\mathbf{d}$ in order to prevent the terms from blowing up, which is similar to the zero-code prior in [27, 20]. In particular, the nonlinear depth map registration attempts to minimize the cost function $\mathcal{C}_{dr}$ defined as
@@
$\hat{z}_{inv_{\ell k}}$ is the estimated metric inverse depth of the $\ell$th feature at time $t_{k}$, and $h_{ds}$ projects the 3D point ${}^{A}\mathbf{p}_{\ell}$ into frame $k$ and extracts the inverse depth. $\sigma_{d}$ is the manually-tuned noise for the depth consistency residual. We also have the zero $\mathbf{c}$, $\mathbf{d}$ prior:
@@ III-C3 Local Mapping Optimization
After registration, as shown in Fig. 5 we perform a sliding-window optimization in the local mapper, with the oldest pose, velocity, and biases fixed. The default sliding window size of our system is 9. Online calibration of the camera intrinsics and extrinsics is also performed in this optimization.
All of the poses are defined in the anchored frame $\{A\}$, which is the gravity-aligned frame centered at the oldest keyframe and is changed after the oldest keyframe is popped from the window. This is similar to robocentric VIO (R-VIO [16]), where the robocentric frame is defined to be the oldest frame in the sliding window, which requires actually estimating the gravity direction. On the other hand, our anchored frame is gravity-aligned – only yaw-rotated when the oldest keyframe is popped off the window – and does not required estimating gravity. While the global pose is not tracked in the local mapper, AB-VINS still calculates it for system output by using the most recent global pose in the memory tree to transform poses in the $\{A\}$ frame into the global. [^3] Note that traversal of the memory tree to calculate the global pose, which is logarithmic in the number of keyframes, is done solely in the mapping thread to keep VIO constant-time.
Besides the standard reprojection error, our local mapping optimization also includes depth consistency cost. Due to having a dense depth map for each keyframe, every bearing in the sliding window is actually its own landmark, and the landmarks simply have to work together with the oldest landmark via the depth consistency cost to keep the system stable. Once the oldest bearing for a particular feature track is older than the sliding window, it becomes a static 3D landmark ${}^{A}\mathbf{p}_{f_{\ell}}$, and replaces the AB feature for the reprojection error and depth consistency. More specifically, the local mapping optimization tries to minimize the follwoing cost function $\mathcal{C}_{lm}$:
@@
where $\hat{z}_{inv_{\ell a}}$ is the metric inverse depth in the anchor frame for this 3D point at keyframe index $a$. Similarly, $\mathbf{r}^{r}_{\ell i}$ is:
@@
IV Mapping with the Memory Tree
After the oldest keyframe is popped off of the local mapping window, it is sent to the mapper thread, which is responsible for finding and solving for loop closures. As shown in Fig. 2, right before entering the mapper, the keyframe receives a global descriptor from NetVLAD [3] and local descriptors from SuperPoint [4], which are typically different from the FAST corners used by VIO. At this point, in order to save RAM memory, all image-resolution dense matrices (e.g., depth map and code mask) are deleted and only sparse portions are maintained at the local descriptor locations. The global descriptor is PCA compressed to 512 dimensions, and searching for loop candidates with this descriptor is extremely efficient. The local descriptors are used to match local keypoints between loop candidates, and if enough matches are found, Horn’s method [45] wrapped in RANSAC is used to compute the relative pose for the pose graph optimization. Upon a loop closure, the loop measurements are sent to the memory tree.
IV-A Tree Structure
The memory tree is a novel data structure proposed to significantly speed up pose graph optimization. As seen in Fig. 6, poses in the memory tree are defined in the parent frame within a binary search tree, while only the root’s pose is in the global frame. As VINS has 4 unobservable directions [46] and drifts in 3D position and yaw, we adopt the 4-DoF poses in the memory tree:
@@ IV-A Tree Structure
where ${}_{i}^{j}\theta_{z}$ is the yaw angle, $\mathbf{R}_{z}$ constructs a yaw-only 3D rotation matrix, and ${}^{i}\mathbf{p}_{j}$ is the gravity-aligned position. The keyframe’s timestamp is the search key for the tree. The memory tree is balanced using the AVL tree rotation rules [47], which guarantee a perfectly balanced tree no matter what. Since the tree is always balanced, calculating the global pose for any keyframe can be done in $\mathcal{O}(\log(N))$ time if there are $N$ nodes in the tree by repeatedly transforming the pose into the same frame as the parent’s pose until the root node is reached. The benefit of defining the poses in this way is that optimizing a single pose also moves the entire subtree underneath the optimized pose’s node.
@@
To limit the memory usage in areas of repeated visitation, redundant nodes are pruned from the memory tree. A node is considered redundant if a certain percent of the map points visible in its keyframe are visible in at least three other keyframes, which is similar to the keyframe culling criteria of ORB-SLAM [6]. After a node is pruned, the tree is rebalanced, again using the AVL balancing rules, to ensure that the height of the tree is $\mathcal{O}(\log(N))$.
IV-B Optimizing Memory Tree
where $\mathcal{S}$ is the set of necessary sequential connections to add (nodes that are next to each other temporally) and $\mathcal{L}$ is the set of necessary loop connections to add. $\bm{\Sigma}_{tree}$ is the manually-tuned diagonal covariance matrix for tree edges, and $\rho_{c}$ is again the Cauchy loss function. Given a 4-DoF relative pose measurement ${}^{i}_{j}\theta_{z}$, ${}^{i}\mathbf{p}_{j}$ between nodes $i$ and $j$, the residual is
where $\{i\}$, $\{j\}$ are the local gravity-aligned frames for nodes $i$ and $j$, and $\{L\}$ is the gravity-aligned frame centered around the lowest common ancestor (LCA) for nodes $i$ and $j$. Finding the LCA can be performed in $\mathcal{O}(\log(N))$ time. Of course the poses for the nodes are defined in the parent frame and not the $\{L\}$ frame, but the poses in the $\{L\}$ frame have to be found by transforming iteratively into the same frame as the parent until the LCA is reached. If optimizing all states in the memory tree, the Hessian structure will be very dense due to correlating more states than usual. Of course, the point of the memory tree is not to optimize every state, but to carefully choose which states to optimize.
Given two loop candidate tree nodes, we can actually move all the nodes needed to close the loop by only optimizing nodes along the shortest path between the two loop nodes. In order to do this, first the LCA must be found between the two. This is because the shortest path between any two nodes in the tree always passes through the LCA. To stabilize the optimization, the root node always remains fixed. Note that this does in most cases allow the first ever pose to move, but we argue that the global pose can be simply taken relative to the first pose. Fig. 7 shows different examples of optimization patterns on the memory tree when optimizing the full path between two loop nodes. Note that edges with only fixed parameters need not be included in the optimization. Thus, not all measurements are added to the optimization problem. The path between the two loop nodes is traversed, and only at most $\mathcal{O}(\log(N))$ sequential relative edges are added, and all loop measurements are checked to see if any of their parameters are included in the set of variable parameters. If so, the loop measurement is added, and if not it is left out of the optimization problem.
While optimizing the full path between two loop nodes is already highly-efficient, an even more efficient solution exists. Since optimizing a single node’s pose moves the entire subtree underneath, it is possible to start with only optimizing the LCA and its two children, and work our way down until the optimization converges in one iteration. We call this the memory tree top-down optimization algorithm. Fig. 8 shows how this works. This solution is not only highly-efficient, but it also solves pose graph SLAM while only relinearizing a constant number of variables. This is because on average only three states are variable (the LCA and its two children), while the rest are fixed. However, in the worst case $\mathcal{O}(\log(N))$ states are set variable.
To robustify against potential bad loop closure measurements (e.g., from visual aliasing), we propose a robust yet efficient $\chi^{2}$ check. Before optimizing the memory tree with a new loop measurement, first the 4-DoF states of all the nodes along the shortest path between the two loop nodes are copied and stored. After optimizing the memory tree, if the new loop residual does not pass a $\chi^{2}$ test with theoretical threshold $\gamma$ (i.e., $||\mathbf{r}^{tree}_{ij}||^{2}_{\bm{\Sigma}_{tree}}\geq\gamma$ if the newest loop measurement is between nodes $i$ and $j$), then the old states are copied back into the memory tree to restore it to what it was before, and the loop measurement is discarded. As such, the proposed AB-VINS is robust to catastrophically bad loop closures while only using $\mathcal{O}(\log(N))$ extra memory.
IV-C Complexity Analysis
@@
Table I shows the complexity analysis of different pose graph optimization algorithms when optimizing with $N$ keyframes. It is evident that our memory tree top-down method is the overall best. Looking at the number of variables affected by optimization (i.e., the number of variables that need to be relinearized), the standard method as in SPA [29] and iSAM [30], with the poses defined in the global frame, affects $\Theta(N)$ variables no matter what. iSAM2 [31] has a chance to affect a constant number of variables under exploration, but typically affects $\mathcal{O}(N)$ variables under a loop closure scenario. The relative method [32] can affect up to $\mathcal{O}(N)$ variables, but on average can affect only a constant number of variables due to only optimizing the beginning and end of each loop. The memory tree top-down method can also affect a constant number of variables, but in the worst case is only $\mathcal{O}(\log(N))$, and only affects a constant number of variables on average even under loop closure scenarios. Moving on to the global pose calculation, defining the poses in the global frame as is standard allows for constant-time global pose calculation. For the relative method, global pose calculation typically takes $\mathcal{O}(N)$ time since each keyframe pose is defined in the previous one’s frame of reference, and thus the global pose is dependent on each previous keyframe. On the other hand, with the memory tree, global pose calculation is typically done in $\mathcal{O}(\log(N))$ time due to the efficient tree structure of the pose definitions. It should be noted that the computation incurred by having to relinearize a larger number of variables is much higher than having to iterate to find the global pose.
Table I also shows the computational complexity of the overall optimization. Note that an average case is not reported here since the complexity depends on the number of loop measurements, which depends on the environment. Thus, there are two worst cases reported for $\mathcal{O}(1)$ and $\mathcal{O}(N)$ loop measurements. With a standard pose graph optimization method, the complexity is linear in the absolute best case, which is the same as having a constant number of loops. In the worst case with $\mathcal{O}(N)$ loops, the complexity is $\mathcal{O}(N^{3})$ [29]. For iSAM2 [31], the optimization can be constant-time in the best case (under exploration) but other than that shares the same complexity with the standard pose graph optimization. With the relative method, the best case arises when the two loop keyframes are very close to each other in time, and there are a constant number of loops. In this case, the method can be constant time. In the worst case with a constant number of loops, the complexity is cubic. This is because $\mathcal{O}(N)$ states can be set variable even with a single loop, and the states are all correlated leading to a dense Hessian structure. Similarly, the complexity of the relative method is cubic with $\mathcal{O}(N)$ loops. Note that this analysis of the relative method assumes that 1) it is applied to pose graph optimization and 2) only the necessary measurements are included in the optimization instead of all the measurements as proposed in the original paper [32]. If all measurements are included in the optimization then the best case becomes $\Omega(N)$ due to having to iterate through all of the sequential edges.
Moving on the the memory tree, with all states set variable, the complexity is cubic in every case. This is due to the cost of LM iterations with a nearly dense Hessian structure. For the full path algorithm, the best case is when there are a constant number of loops and the two loop nodes are near the root of the tree. In this case, the complexity is constant. In the worst case with a constant number of loops, the complexity is $\mathcal{O}((\log(N))^{3})$, which is sublinear. This is due to the cost of LM iterations with $\mathcal{O}(\log(N))$ states that are all correlated. In the worst case with $\mathcal{O}(N)$ loops, the complexity is $\mathcal{O}(N\log(N))$ due to having to find the LCA for $\mathcal{O}(N)$ loops to see which measurements should be included in the optimization (which is typically more efficient than the LM iterations). For the top-down method, the complexity is the same as the full path algorithm since $\mathcal{O}(\log(N))$ states can be variable in the worst case. If only a constant number of states are variable (which is the average case) with a constant number of loops, the complexity is $\mathcal{O}((\log(N))^{2})$ due to having to find the LCA for the $\mathcal{O}(\log(N))$ relative edges. Since AB-VINS prunes many duplicate nodes, loop measurements are typically discarded, leading to a constant number of loop measurements. While AB-VINS allows for at most $\mathcal{O}(N)$ loop measurements, some systems such as ORB-SLAM3 [6] allow for up to $\mathcal{O}(N^{2})$ loops. If this is the case, the worst case for the full path and top-down algorithms becomes $\mathcal{O}(N^{2}\log(N))$, which is still better than the cubic complexity of other methods.
V Experimental Results
We implement the proposed AB-VINS in an efficient C++ codebase. Ceres solver [48] with single-thread execution is used for all optimizations. Automatic differentiation provided by ceres::Jet is used in all factors except for preintegration and the reprojection error from a static point used in tracking to static map, which use analytical Jacobians. Unlike numerical differentiation via finite differences, automatic differentiation does not come with a large computational overhead or loss of accuracy. We found automatic differentiation to be especially useful for developing the memory tree relative pose factor, since the analytical Jacobians for it are complicated. A desktop with an i5-6600K CPU and RTX 2070 Super GPU with 16GB of RAM is used in all experiments.